A solenoid with a cross-sectional area of 1.16×10−3m2 is 0.770 m long and has 900 turns per meter.

Part A

Find the induced emf in this solenoid if the current in it is increased from 0 to 2.50 A in 31.8 ms

__________mV
Answer #1

Part A:

Given data,

l = 0.7 m

N = 900 turns

dl = 2.50 A

t = 0.0318 sec

A = 1.16×10−3m2

We know that,

EmF = (B * A * n) /t Where l is length , t is time , dl is change in current

B = u0 * N * I

EmF induced = (u0 * N)* N * l* A * dI/dt

= 1.26 * 10^-6 * 900 * 900 * 0.7 * 1.16 *10-3 * 2.50 / 0.0318

= 0.002071818 / 0.0318

= 0.0651

= 65.1 x 10-3 V

= 65.1 mV